Find the inverse function for the function f(x) = x^(1/3) - 3
I'm having trouble understanding how to do problems like these. If anyone could show me, I'd appreciate it. Thanks.

Answer by funmath(2873)

Find the inverse function for the function f(x) = x^(1/3) - 3
Let f(x)=y
y=x^(1/3)-3 Let x's be y's and y's be x's and solve for y
x=y^(1/3)-3
x+3=y^(1/3)-3+3
x+3=y^(1/3)
(x+3)^3=(y^(1/3))^3
(x+3)^3=y let y be f^-1(x)
f^-1(x)=(x+3)^3
Happy Calculating!!!